OpenStudy (jeanfifa09):

Find the nth term of the sequence 1,-4,9,-16,... A) an=32 B) an=-2n C) an=(-2)^n D) an=(-1)^n • (n)^2 E) an=n^2

OpenStudy (jeanfifa09):

I need help please . I can't match any of the answers

OpenStudy (welshfella):

hint : all the numbers are squares except there is a negative sign in front of the even number squares

OpenStudy (welshfella):

- so there is an n^2 involved in the nth term

OpenStudy (jeanfifa09):

Ok sorry why is there a negative only on the even numbers ,what formula match the sequence

OpenStudy (welshfella):

yes - i was thinking that an was = D but thats not correct

OpenStudy (welshfella):

because first term when n = 1 would be (-1)^1 * (1)^2 = -1 * 1 = -1 not 1

OpenStudy (welshfella):

None of the options fit the sequence correct an would be (-1)^(n-1) * (n)^2 here a1 = (-1)^0 * (1)^2 = 1*1 = 1 a2 = (-1)^1 * (2)^2 = -1*4 = -1 a3 = (-1)^2 * (3)^2 = 1 * 9 = 9

OpenStudy (jeanfifa09):

I know ,so all of them are wrong ?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

could there be a typo somewhere?

OpenStudy (welshfella):

D is the closest

OpenStudy (welshfella):

D would give a sequence -1,4,-9,16

OpenStudy (jeanfifa09):

Oh ok ,thank you

OpenStudy (welshfella):

yw